what is the orbit radius of a 300kg satellite that completes one circular orbit around the earth every 3.6hr?


1
Expert's answer
2022-01-16T13:09:08-0500

Answer


Time period


T=3.6hr=12900 sec.


Mass M=300kg


Using kepler third law

"T^2=\frac{4\pi^2}{GM}r^3"

"r^3=\frac{T^2GM}{4\pi ^2}\\"

"\frac{(12960)^2\times 6.67\times10^{-11}\times300}{4\times(3.14)^2}\\=\frac{3.361}{39.44}\\=0.085"

So this is radius of orbit.